# Incentives in Federated Learning with Heterogeneous Agents

### v1

@@ Abstract
Federated learning promises significant sample-efficiency gains by pooling data across multiple agents, yet incentive misalignment is an obstacle: each update is costly to the contributor but boosts every participant. We introduce a game-theoretic framework that captures heterogeneous data: an agent’s utility depends on who supplies each sample, not just how many. Agents aim to meet a PAC-style accuracy threshold at minimal personal cost. We show that uncoordinated play yields pathologies: pure equilibria may not exist, and the best equilibrium can be arbitrarily more costly than cooperation. To steer collaboration, we analyze the cost-minimizing contribution vector, prove that computing it is NP-hard, and derive a polynomial-time linear program that achieves a logarithmic approximation. Finally, pairing the LP with a simple pay-what-you-contribute rule—each agent receives a payment equal to its sample cost—yields a mechanism that is strategyproof and, within the class of contribution-based transfers, is unique.
@@ 1 Introduction
Federated learning (FL) is a collaborative training framework in which multiple agents—each holding a distinct dataset—jointly optimize a global model while keeping data local. Collaboration allows agents to tap into information spread across heterogeneous records—for example, a network of hospitals pooling imaging data from distinct patient demographics to detect rare conditions sooner. Although each agent could independently train a model, collaboration offers higher accuracy or comparable performance with significantly fewer examples [1], enhancing both individual and federation-wide welfare.
However, realizing these gains hinges on incentives. Contributing model updates incurs costs in compute, bandwidth, curation effort, and privacy risk, while the global model produced by collective learning is a non-excludable public good: once trained, every agent benefits from its accuracy regardless of individual effort. This asymmetry invites a classic free-rider dilemma [2, 3, 4]: as one agent’s data lifts others’ accuracy while the contributor alone incurs the cost, each participant is tempted to trim its share once its own target is met. The resulting free-riding slows training and can ultimately exhaust the data pool that makes FL viable.
Consequently, federated learning can be viewed as a strategic game: each agent picks a contribution level to maximize a private utility that trades its own labeling cost against the benefit of the joint model. Existing models of incentives in federated learning either assume each agent’s utility depends on the local labeled data distributions of all agents [5], or treat agents as homogeneous, which crucially implies that data are exchangeable and each agent’s utility depends solely on the total amount of data contributed across all agents [4, 6].
However, in typical federated learning scenarios, agents face heterogeneous data distributions and are primarily interested in improving performance on their own local distribution. This phenomenon—local data typically yielding greater marginal utility for local performance than data from other agents—has been well documented in the personalization and domain adaptation literature [7, 8, 9], and we empirically confirm that it is present in federated learning, as shown in Figure˜1. We therefore model utilities that depend on who supplies each sample, not just how many.
@@
Figure 1: An agent’s expected loss falls as a larger share of a fixed training set comes from their own distribution. With a budget of $m$ data points, we sample $\lambda m$ from Agent 1 and $(1-\lambda)m$ from Agent 2 on FEMNIST [10], train a classifier, and repeat 100100 times. For each $m$, Agent 1’s loss decreases monotonically in $\lambda$, confirming that data are not exchangeable—utilities depend on who contributes. Full details in Section˜A.1.
@@
Early work in FL centered on communication-efficient optimization and fairness under fully cooperative, i.i.d. or average-loss assumptions, with no pricing of data or modeling of strategic behavior. For instance, FedAvg optimizes average loss, while Agnostic FL and q-FFL reweight loss for worst-case or fairness gains, all assuming truthful reports [11, 12, 13]. Blum et al. [1] introduced a collaborative PAC model and showed that pooling across $k$ heterogeneous tasks cuts sample complexity by $\widetilde{O}(\log k)$, but participation is mandatory and incentives are ignored.
Subsequent work introduced incentives but frequently assumed that agents’ data are exchangeable, so that model accuracy depends only on the total sample count contributed across all participants. Under this framework, various incentive strategies—such as transfer payments or reputational rewards—have been proposed to combat free-riding. For example,  Karimireddy et al. [4] tie each agent’s model quality to its data contribution (no external payments), whereas  Murhekar et al. [6] design budget-balanced monetary transfers that implement welfare-maximizing equilibria; and [14, 15, 16, 17, 18] rely on reputation mechanisms and credit sharing. These designs treat data as interchangeable: model quality depends solely on the total sample count, so the marginal value of each sample ignores who provides it, overlooking realistic heterogeneity across data distributions.
A separate line of work studies heterogeneous data, aiming to capture how collaboration might form among agents with distinct interests.  Donahue and Kleinberg [19], and  Hasan [20] analyze coalition and Nash stability in model-sharing games and show that agents split into sub-coalitions when a global model biases some distributions.  Blum et al. [5] further show that in a personalized PAC game with distribution-specific payoffs, pure Nash equilibria may not exist, and when they do, they can be arbitrarily inefficient, underscoring the fragility of cooperation absent incentives. These game-theoretic models emphasize that heterogeneity can severely complicate collaboration: individual incentives may fail to align with socially optimal pooling of data. Our work extends these papers by introducing a concrete utility model based on PAC-style threshold guarantees where each agent requires that the global model meet a distribution-specific accuracy threshold with high confidence.
Recent studies tackle truthfulness under heterogeneity: peer-prediction payments [21], budget-balanced truthful gradient schemes [22], and an optimal truthful mechanism for data sharing with interdependent valuations [23]; yet none compute minimum-cost allocations that meet each agent’s welfare target.
@@ 2.1 Setup and Learning Protocol
where $\text{err}_{S}(h)$ is the empirical error of hypothesis $h$ on dataset $S$. For any hypothesis $h$, marginal data distribution $\mathcal{D}$, and target function $h^{\star}$, the generalization error is defined as
@@
Stages 1 and 2 define our collaborative game. Each agent $i$ chooses an integer contribution level $m_{i}$. The server draws $m_{i}$ i.i.d. samples from $\mathcal{D}_{i}$ for each agent, pools the labeled data into $S$, and trains a single global ERM predictor $\widehat{h}(S)$ over $\mathcal{H}$. The next subsection specifies the payoffs from this global model and the cost of contributing data.
@@ 2.2 PAC Accuracy Objective
This self-sufficiency implies that failing to meet the accuracy threshold results in strictly lower utility than meeting it (since an agent can always fall back on solo training to obtain the benefit, albeit at potentially high cost).
@@ 3 The Inefficiency of Equilibria
This example shows that some equilibria can be much costlier than the social optimum even though a cost-minimal equilibrium exists. How large can the best equilibrium’s cost be? We formalize it with the Price of Stability (PoS) [24]: the ratio of total samples in the least-cost equilibrium to those in the planner’s optimum. PoS captures the cost of decentralization. Even the best self-enforcing outcome can require far more samples than the cooperative minimum. In the following, we show that PoS can be arbitrarily large:
@@ 4.2 Approximation via Linear Programming
For any ordered pair $(h_{1},h_{2})$ satisfying this condition, the probability that no sample lands in the set $\{x:h_{1}(x)\neq h_{2}(x)\}$ is $\prod_{i\in\mathcal{A}}\bigl(1-\mathcal{D}_{i}(\{h_{1}\neq h_{2}\})\bigr)^{m_{i}},$ which is log-linear in $\mathbf{m}$. We can therefore convert Equation˜4 into an LP with polynomially many constraints, which can be solved efficiently. As an empirical check, Section˜A.2 validates this LP allocation on a finite hypothesis class by comparing to the true optimum $\sum_{i}m_{i}^{\star}$ across varying $|\mathcal{H}|$.
@@ 5 Mechanism Design with Approximate Solutions
In Section˜4 we studied the planner’s computational problem under full information and control. Here we keep prescriptive control but drop full information and reintroduce strategy under coordination. In the uncoordinated game of Section˜3 agents act by choosing their own sample sizes. With the planner fixing sample counts, agents can only influence outcomes by misreporting their local distributions, which shifts the LP constraints from Section˜4, [^4] The information that agents must provide to the server is finite and explicit. Our LP needs, for each pair $h_{1},h_{2}\in\mathcal{H}$, the probability that agent $i$’s data lie in the disagreement set $\{x:h_{1}(x)\neq h_{2}(x)\}$. If $\mathcal{H}$ is finite, this is a vector with at most $|\mathcal{H}|^{2}$ entries. If $\mathcal{H}$ is infinite, Section D.2 replaces $\mathcal{H}$ with a finite $\gamma$-cover whose size is polynomial in $d=\operatorname{VCdim}(\mathcal{H})$. Agents then report disagreement masses only for pairs in the cover, so the information burden remains finite and explicit. and thus the computed contribution vector. [^5] Enforcing prescribed contributions without accessing raw data raises a verification challenge. Practical approaches include cryptographic proofs and auditable contribution records [25, 26]. Building on that LP, in this section, we design a mechanism that is strategyproof and unique within a broad contribution-based class, blocking this manipulation and addressing the inefficiency in Section˜3. Together, these results highlight the contrast between uncoordinated and coordinated settings.
Specifically, we now regard the agents’ local marginal data distributions as private information that must be reported to the mechanism. To alleviate the resulting incentive issues, we allow payments to the agents. The sequence of events, therefore, is as follows:
@@ 5.2 Alternatives to Pay What You Contribute?
where $q_{i}(\bm{\mathcal{D}}_{-i})$ is a term independent of the report $\mathcal{D}_{i}$ of agent $i$. By standard arguments [27], the VCG mechanism ensures that each agent reports truthfully.
An obstacle to using VCG directly is that computing the optimal contribution vector is computationally hard (Theorem˜2). In the algorithmic mechanism design [28] literature, however, there are various mechanisms that overcome computational hardness by augmenting approximation algorithms with clever payment schemes, including ones inspired by VCG [29, 30]. Is there such a rich mechanism design space in our setting? We give a partial negative answer to this question, showing that the PWYC mechanism is, in a qualified sense, unique. [^6] This uniqueness is unfortunate, since a budget-balanced mechanism would be preferable. Even in standard mechanism design settings, particularly the celebrated VCG mechanism, exact budget balance is generally impossible, so the planner must make net payments to align incentives. Still, PWYC need not put all costs on the planner. By choosing constants $C_{i}$ in Equation 5, the planner can recover part of the cost while preserving strategyproofness.
@@ Appendix A Empirical Validation and Illustrative Simulations
We build on the LEAF benchmark [10], which provides realistic federated datasets naturally partitioned by user. We selected two datasets, FEMNIST and Shakespeare, to cover both vision and textual tasks. These specific datasets were chosen because they have the highest average number of datapoints per user among the LEAF collection [10]. FEMNIST, built on Extended MNIST [31, 32], consists of handwritten character images partitioned by writer. Shakespeare is constructed from The Complete Works of William Shakespeare [33, 11], partitioned by speaking role, with each role in each play treated as a distinct agent.

### v2

@@ Abstract
Federated learning promises significant sample-efficiency gains by pooling data across multiple agents, yet incentive misalignment is an obstacle: each update is costly to the contributor but boosts every participant. We introduce a game-theoretic framework that captures heterogeneous data: an agent’s utility depends on who supplies each sample, not just how many. Agents aim to meet a PAC-style accuracy threshold at minimal personal cost. We show that uncoordinated play yields pathologies: pure equilibria may not exist, and the best equilibrium can be arbitrarily more costly than cooperation. To steer collaboration, we analyze the cost-minimizing contribution vector, prove that computing it is NP-hard, and derive a polynomial-time linear program that achieves a logarithmic approximation. Finally, pairing the LP with a simple pay-what-you-contribute rule—each agent receives a payment equal to its sample cost—yields a mechanism that is strategy-proof and, within the class of contribution-based transfers, is unique.
@@ 1 Introduction
Federated learning (FL) is a collaborative training framework in which multiple agents—each holding a distinct dataset—jointly optimize a global model while keeping data local. Collaboration allows agents to tap into information spread across heterogeneous records—for example, a network of hospitals pooling imaging data from distinct patient demographics to detect rare conditions sooner. Although each agent could independently train a model, collaboration offers higher accuracy or comparable performance with significantly fewer examples (Blum et al., 2017), enhancing both individual and federation-wide welfare.
However, realizing these gains hinges on incentives. Contributing model updates incurs costs in compute, bandwidth, curation effort, and privacy risk, while the global model produced by collective learning is a non-excludable public good: once trained, every agent benefits from its accuracy regardless of individual effort. This asymmetry invites a classic free-rider dilemma (Yang et al., 2015; Ahmed and Choi, 2023; Karimireddy et al., 2022): as one agent’s data lifts others’ accuracy while the contributor alone incurs the cost, each participant is tempted to trim its share once its own target is met. The resulting free-riding slows training and can ultimately exhaust the data pool that makes FL viable.
Consequently, federated learning can be viewed as a strategic game: each agent picks a contribution level to maximize a private utility that trades its own labeling cost against the benefit of the joint model. Existing models of incentives in federated learning either assume each agent’s utility depends on the local labeled data distributions of all agents (Blum et al., 2021), or treat agents as homogeneous, which crucially implies that data are exchangeable and each agent’s utility depends solely on the total amount of data contributed across all agents (Karimireddy et al., 2022; Murhekar et al., 2024).
However, in typical federated learning scenarios, agents face heterogeneous data distributions and are primarily interested in improving performance on their own local distribution. This phenomenon—local data typically yielding greater marginal utility for local performance than data from other agents—has been well documented in the personalization and domain adaptation literature (Ben-David et al., 2006; Bhunia et al., 2021; Hsu et al., 2019), and we empirically confirm that it is present in federated learning, as shown in Figure˜1. We therefore model utilities that depend on who supplies each sample, not just how many.
@@
Figure 1: An agent’s expected loss falls as a larger share of a fixed training set comes from their own distribution. With a budget of $m$ data points, we sample $\lambda m$ from Agent 1 and $(1-\lambda)m$ from Agent 2 on FEMNIST (Caldas et al., 2018), train a classifier, and repeat 100100 times. For each $m$, Agent 1’s loss decreases monotonically in $\lambda$, confirming that data are not exchangeable—utilities depend on who contributes. Full details in Section˜A.1.
@@
Early work in FL centered on communication-efficient optimization and fairness under fully cooperative, i.i.d. or average-loss assumptions, with no pricing of data or modeling of strategic behavior. For instance, FedAvg optimizes average loss, while Agnostic FL and q-FFL reweight loss for worst-case or fairness gains, all assuming truthful reports (McMahan et al., 2017; Mohri et al., 2019; Li et al., 2020). Blum et al. (2017) introduced a collaborative PAC model and showed that pooling across $k$ heterogeneous tasks cuts sample complexity by $\widetilde{O}(\log k)$, but participation is mandatory and incentives are ignored.
Subsequent work introduced incentives but frequently assumed that agents’ data are exchangeable, so that model accuracy depends only on the total sample count contributed across all participants. Under this framework, various incentive strategies—such as transfer payments or reputational rewards—have been proposed to combat free-riding. For example,  Karimireddy et al. (2022) tie each agent’s model quality to its data contribution (no external payments), whereas  Murhekar et al. (2024) design budget-balanced monetary transfers that implement welfare-maximizing equilibria; and (Lin et al., 2019; Kang et al., 2019; Sarikaya and Ercetin, 2019; Ding et al., 2020; Fraboni et al., 2021) rely on reputation mechanisms and credit sharing. These designs treat data as interchangeable: model quality depends solely on the total sample count, so the marginal value of each sample ignores who provides it, overlooking realistic heterogeneity across data distributions.
A separate line of work studies heterogeneous data, aiming to capture how collaboration might form among agents with distinct interests.  Donahue and Kleinberg (2021), and  Hasan (2021) analyze coalition and Nash stability in model-sharing games and show that agents split into sub-coalitions when a global model biases some distributions.  Blum et al. (2021) further show that in a personalized PAC game with distribution-specific payoffs, pure Nash equilibria may not exist, and when they do, they can be arbitrarily inefficient, underscoring the fragility of cooperation absent incentives. These game-theoretic models emphasize that heterogeneity can severely complicate collaboration: individual incentives may fail to align with socially optimal pooling of data. Our work extends these papers by introducing a concrete utility model based on PAC-style threshold guarantees where each agent requires that the global model meet a distribution-specific accuracy threshold with high confidence.
Recent studies tackle truthfulness under heterogeneity: peer-prediction payments (Pang et al., 2022), budget-balanced truthful gradient schemes (Chakarov et al., 2024), and an optimal truthful mechanism for data sharing with interdependent valuations (Chen et al., 2022); yet none compute minimum-cost allocations that meet each agent’s welfare target.
@@ 2.1 Setup and Learning Protocol
where $\text{err}_{S}(h)$ is the empirical error of hypothesis $h$ on dataset SS. [^1] In practice, algorithms such as FedAvg (McMahan et al., 2017) only approximate the pooled-ERM solution through local updates and model aggregation, often without accessing raw examples, which remain on users’ devices and are never aggregated on the server. In our work, the optimal ERM solution is used as an upper-bound reference point, which lets us abstract away optimization dynamics and focus on incentives and sample complexity. See Section 6 for more discussion. For any hypothesis $h$, marginal data distribution $\mathcal{D}$, and target function $h^{\star}$, the generalization error is defined as
@@
Stages 1 and 2 define our collaborative game. Each agent $i$ chooses an integer contribution level $m_{i}$. The server draws $m_{i}$ i.i.d. samples from $\mathcal{D}_{i}$ for each agent, pools the labeled data into $S$, and trains a single global ERM predictor $\widehat{h}(S)$ over $\mathcal{H}$. The next subsection, Section˜2.2, specifies the payoffs from this global model and the cost of contributing data.
@@ 2.2 PAC Accuracy Objective
This self-sufficiency implies that failing to meet the accuracy threshold results in strictly lower utility than meeting it (since an agent can always fall back on solo training to obtain the benefit, albeit at potentially high cost). Therefore, the agent-level decision problem in Equation˜3 is to choose $m_{i}$ so as to minimize its own cost $c_{i}m_{i}$ subject to satisfying its $(\varepsilon,\delta)$-accuracy requirement in Equation˜2.
This binary threshold payoff can represent a broad class of quasi-linear utility functions, and captures richer behavior than its simple indicator form suggests at first glance. Under regularity assumptions, Appendix˜B shows via a Lagrange-multiplier argument that any optimal solution to above constrained problem is also optimal for some unconstrained objective that trades off continuous accuracy against cost; in particular, there is a multiplier $\lambda_{i}$ such that the optimal contribution $m_{i}^{*}$ that meets the PAC constraint also maximizes a quasi-linear utility that penalizes cost at rate $\lambda_{i}$. Intuitively, $\lambda_{i}$ is the agent’s shadow price for accuracy, and the point where our indicator flips from 0 to $1$ is exactly where the marginal benefit of additional accuracy falls to this price. Thus the threshold model is a concise way to capture the saturation behavior of a wide family of smooth, monotone, diminishing-returns utilities, rather than an assumption that agents literally have zero value below the target.
@@ 3 The Inefficiency of Equilibria
This example shows that some equilibria can be much costlier than the social optimum even though a cost-minimal equilibrium exists. How large can the best equilibrium’s cost be? We formalize it with the Price of Stability (PoS) (Anshelevich et al., 2008): the ratio of total samples in the least-cost equilibrium to those in the planner’s optimum. PoS captures the cost of decentralization. Even the best self-enforcing outcome can require far more samples than the cooperative minimum. In the following, we show that PoS can be arbitrarily large:
@@ 4.2 Approximation via Linear Programming
For any ordered pair $(h_{1},h_{2})$ satisfying this condition, the probability that no sample lands in the set $\{x:h_{1}(x)\neq h_{2}(x)\}$ is $\prod_{i\in\mathcal{A}}\bigl(1-\mathcal{D}_{i}(\{h_{1}\neq h_{2}\})\bigr)^{m_{i}},$ which is log-linear in $\mathbf{m}$. We can therefore convert Equation˜4 into an LP with polynomially many constraints, which can be solved efficiently.
@@
Any server that implements this LP to coordinate contributions, as in Section˜5, interacts with agents only through the information the LP requires, and two practical aspects are worth noting. First, the LP needs from each agent only a finite, explicitly specified vector probabilities, so the communication burden is well defined and finite. [^4] Formally, for each pair $h_{1},h_{2}\in\mathcal{H}$ the planner needs, from each agent $i$, the probability that its data lie in the disagreement set $\{x:h_{1}(x)\neq h_{2}(x)\}$, that is, the quantity $\mathcal{D}_{i}(\{x:h_{1}(x)\neq h_{2}(x)\})$. If $\mathcal{H}$ is finite, this yields at most $|\mathcal{H}|^{2}$ numbers per agent. If $\mathcal{H}$ is infinite, Section D.2 replaces $\mathcal{H}$ with a finite $\gamma$-cover $\overline{\mathcal{H}}$ whose size is polynomial in $d=\operatorname{VCdim}(\mathcal{H})$. These are summary statistics that can be estimated locally or on a shared unlabeled validation set, rather than raw examples. Second, these quantities are summary distributional statistics that each agent can compute locally from its own data. Sharing only these statistics, rather than raw examples, ensures the server never receives individual training examples.
Formally, the next theorem summarizes the approximation guarantee of this LP relaxation; the full proof appears in Section˜D.2. As an empirical check, Section˜A.2 validates this LP allocation on a finite hypothesis class by comparing to the true optimum $\sum_{i}m_{i}^{\star}$ across varying $|\mathcal{H}|$.
@@ 5 Mechanism Design with Approximate Solutions
In Section˜4 we studied the planner’s computational problem under full information and control. Here we keep prescriptive control but drop full information and reintroduce strategy under coordination. In the uncoordinated game of Section˜3 agents act by choosing their own sample sizes. With the planner fixing sample counts, agents can only influence outcomes by misreporting their local distributions, which shifts the LP constraints from Section˜4, and thus the computed contribution vector. [^5] Enforcing prescribed contributions without accessing raw data raises a verification challenge. Practical approaches include cryptographic proofs and auditable contribution records (Ma et al., 2024; Kumar et al., 2025). Building on that LP, in this section, we design a mechanism that is strategyproof and unique within a broad contribution-based class, blocking this manipulation and addressing the inefficiency in Section˜3. Together, these results highlight the contrast between uncoordinated and coordinated settings.
Specifically, we now regard the agents’ local marginal data distributions as private information that must be reported to the mechanism. Their per-sample costs $c_{i}$ remain known parameters. [^6] For many applications, it is natural to treat per-sample costs as known or contractually specified: platform may post a fixed price per labeled example, meter and bill bandwidth or compute usage, or sign annotation contracts that set a per-label rate up front, so that $c_{i}$ is a parameter of the environment rather than private strategic information. To alleviate the resulting incentive issues, we allow payments to the agents. The sequence of events, therefore, is as follows:
@@ 5.2 Alternatives to Pay What You Contribute?
where $q_{i}(\bm{\mathcal{D}}_{-i})$ is a term independent of the report $\mathcal{D}_{i}$ of agent $i$. By standard arguments (Nisan, 2007), the VCG mechanism ensures that each agent reports truthfully.
An obstacle to using VCG directly is that computing the optimal contribution vector is computationally hard (Theorem˜2). In the algorithmic mechanism design (Nisan and Ronen, 2001) literature, however, there are various mechanisms that overcome computational hardness by augmenting approximation algorithms with clever payment schemes, including ones inspired by VCG (Lehmann et al., 2002; Dobzinski, 2007). Is there such a rich mechanism design space in our setting? We give a partial negative answer to this question, showing that the PWYC mechanism is, in a qualified sense, unique.
This uniqueness is unfortunate, since the absence of a budget-balanced mechanism naturally raises the concern that shifting costs to the planner could, in a sense, diminish some of the benefits of FL training in the first place. We view the uniqueness of PWYC as a substantive, if somewhat negative, insight: if one insists on dominant-strategy truthfulness and approximate efficiency in a heterogeneous FL game, then some external budget or cost sharing is unavoidable, much like in other public-goods settings where reimbursement is often needed to align incentives. Even in standard mechanism design settings, including the celebrated VCG mechanism, exact budget balance is generally impossible, so the planner must make net payments to align incentives. Importantly, PWYC does not force the planner to shoulder all costs: the constants $C_{i}$ in Equation˜5 can be set to recover part of the spend from participants or downstream users while preserving strategyproofness.
@@ 6 Discussion
Centralized ERM as a benchmark.
In this work, the ERM solution on the pooled dataset $S$ is used only as a theoretical benchmark, not as an implementation assumption. The raison d’être of FL is to avoid aggregating raw data on a central server, so directly minimizing the joint empirical loss over all local data imagines having everything in one place. In practice, algorithms such as FedAvg approximate the pooled-ERM solution via local updates and model aggregation, without ever materializing a pooled dataset. In our analysis, this centralized ERM solution serves as the reference point: it defines the PAC constraints and certifies that, if one were to train on the pooled draw of size $m$, the resulting model would satisfy each agent’s $(\varepsilon_{i},\delta)$ requirement. This abstraction lets us ignore optimization dynamics and focus on incentives and sample complexity, while any FL protocol that gets sufficiently close to this benchmark inherits the same qualitative guarantees. Using centralized ERM as an upper-bound reference is standard in FL theory (Mohri et al., 2019; Li et al., 2020; Blum et al., 2021).
@@ Appendix A Empirical Validation and Illustrative Simulations
We build on the LEAF benchmark (Caldas et al., 2018), which provides realistic federated datasets naturally partitioned by user. We selected two datasets, FEMNIST and Shakespeare, to cover both vision and textual tasks. These specific datasets were chosen because they have the highest average number of datapoints per user among the LEAF collection (Caldas et al., 2018). FEMNIST, built on Extended MNIST (LeCun, 1998; Cohen et al., 2017), consists of handwritten character images partitioned by writer. Shakespeare is constructed from The Complete Works of William Shakespeare (Shakespeare and others, 1989; McMahan et al., 2017), partitioned by speaking role, with each role in each play treated as a distinct agent.